40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Jul 19 04:44:09 2018, Annalisa, Terra, Summary, Thermal Compensation, frequency shift observed with heating! scanning_fine_2018-07-19-07-32-08_parsed.pdfscanning_fine_2018-07-19-06-57-47_parsed.pdfYscan_scanning_parsed_2am.txt.pdf
    Reply  Fri Jul 20 07:43:54 2018, Steve, Summary, SUS, ETMY  ETMY_leveling.pngETMY.png
    Reply  Sat Jul 21 01:06:49 2018, gautam, Summary, Thermal Compensation, Y arm locking 
       Reply  Sat Jul 21 14:03:19 2018, Koji, Summary, Thermal Compensation, Y arm locking 
Message ID: 14094     Entry time: Sat Jul 21 01:06:49 2018     In reply to: 14086     Reply to this: 14096
Author: gautam 
Type: Summary 
Category: Thermal Compensation 
Subject: Y arm locking 

I implemented this today. For now, the LSC output matrix is set to actuate on MC2 for Y arm locking. As expected, the transmission was much more stable, and the PLL control signal RMS was also reduced by factor of ~3. MC2 control signal does pick up a large (~2000 cts) DC component over a few hours, so we need to relieve this periodically.

Now that we have a workable ASS for the Y arm as well, we should be able to have more confidence in returning to the same beam spot position on the ETM and staying there during a scan using this technique.

The main improvement to be trialled next in the scanning is to improve the speed of scanning. As things stand, my script takes ~2.5 seconds per datapoint. If we can cut this in half, that'd be huge. On Wednesday night, we were extraordinarily lucky to avoid MC3 glitching, EPICS/slow machine failures, and GPIB freezes. Today, the latter reared its head. Fortunately, since I'm dumping data to file for each datapoint, this means we at least have data till the GPIB freeze.

Quote:

For future measurements, we should consider locking the IMC length to the arm cavity - this would eliminate such alignment drifts, and maybe also make the PLL control signal RMS smaller. 


Not related to this work: Terra, Sandrine, Keerthana and I cleaned up the lab a bit today, and made better cable labels. Aaron and I have to clean up the OMC area a bit. Huge thanks to Steve for taking care of our mess elsewhere in the lab!

ELOG V3.1.3-